OpenStudy (anonymous):

solve- y-5<6 y<1 y<11 y<-11 y<30

OpenStudy (anonymous):

this _ should be under the sign <.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

Are those the answer choices to the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So add 5 to both sides

OpenStudy (anonymous):

And you get y<11

OpenStudy (anonymous):

So it would be answer choice B

OpenStudy (anonymous):

8>d-2 d<10 d>10 d<6 d>4

jhonyy9 (jhonyy9):

d< 8+2

OpenStudy (znappydooz):

From wat i can tell, it equals y<-11

OpenStudy (znappydooz):

Lol check it out urself if u dont believe me :P I got the answer on mathway

OpenStudy (znappydooz):

Here's the link: https://www.mathway.com/ it really helped me when I was in algebra

OpenStudy (anonymous):

No it would not be a negative eleven, its a positive eleven

OpenStudy (znappydooz):

I'm not so good in algebra so i use Mathway :P it gives me the right answer most of the time. Go check it out for urself and see if it's right cuzz maybe its wrong, who knows

OpenStudy (anonymous):

No, I solved it by hand. Sometimes websites are not useful. I'll plug in 11 to the equation and see what happens. will change the symbol, to a "=" sign. y-5=6. 11-5=6, and 6=6, correct? now lets see -11, -11-5=6. -16 does NOT equal 6. So, it is positive 11.

OpenStudy (znappydooz):

Oh... I see, I think you might be right

OpenStudy (znappydooz):

You are right ^_^ My bad! Good job, again, I'm bad with this stuff. Medal for you lol ;) It is y<11